# Coupling between small polarons and ferroelectricity in BaTiO3

### v3

@@ Abstract
In this study, we investigate the formation of electron and hole small polarons in the prototypical ferroelectric material BaTiO3, with a focus on their interaction with ferroelectric distortive fields. To accurately describe the ferroelectric phase in electronically correlated BaTiO3, we employ the HSE06 hybrid functional, which addresses the limitations of conventional DFT and DFT+U models, providing a more precise depiction of both ferroelectric and polaronic behaviors. Our analysis spans three structural phases of BaTiO3: cubic, tetragonal, and rhombohedral. We uncover a unique phase-dependent trend in electron polaron stability, which progressively increases across the structural phases, peaking in the rhombohedral phase due to the constructive coupling between the polaron and ferroelectric phonon fields. In contrast, hole polarons exhibit a stability pattern largely unaffected by the phase transitions. Furthermore, we observe that polaron self-trapping significantly alters the local ferroelectric distortive pattern, which propagates to neighboring sites but has a minimal effect on the long-range macroscopic spontaneous polarization. Charge trapping is also associated with localized spin formation, opening new possibilities for enhanced functionalities in multiferroic materials.
@@ I Introduction
Figure 1: (a) Cubic BaTiO3 with zero Ti off-centering ($>$393K) (b) Tetragonal BaTiO3 with Ti off-centering in [001] direction (394K - 278K) and (c) Rhombohedral BaTiO3 with Ti off-centering in the [111] direction($<$183K). The insets represent the Ti-O bond-lengths in each phase.
@@
Marked by a spontaneous electric polarization that can be reversed by an external electric field, ferroelectric (FE) materials are vital in a wide range of technological applications [1, 2, 3, 4]. Capitalizing on this unique switchable polarization property, their applications range from capacitors and transducers to non-volatile memory devices [5, 6, 7]. Due to its impressive FE properties, Barium titanate (BaTiO3) has been the subject of exhaustive research [8, 9]. This also includes significant attention to understanding the impact of localized states, such as those induced by defects, on its FE behaviour. BaTiO3 is particularly interesting because it undergoes multiple phase transitions from its high symmetric and high-temperature cubic phase to the room-temperature tetragonal phase (between 394 K and 278 K), further to the orthorhombic phase (between 278 K and 183 K), and finally to the low-temperature rhombohedral phase (below 183 K). Studying the cubic, tetragonal, and rhombohedral phases opens up an avenue for investigating the interaction of localized states with the structure and electronic states across these temperature-dependent phases.
FE behaviour and dielectric properties are altered due to the presence of defects that introduce localized electronic states within the band structure. In perovskites such as BaTiO3 and ZrPtO4, the presence of oxygen vacancies act as electron donors, potentially creating defect dipoles that pin domain walls and thereby impacting polarization switching [10, 11, 12, 13]. Other studies show that ferroelectricity decreases due to cation vacancies, such as titanium vacancies in BaTiO3 [14, 15]. Similarly, for SrTiO3, intrinsic defects like antisite titanium induce local polarization and lead to the formation of polar nanoregions [16, 17, 18]. Furthermore, studies also reveal that the introduction of foreign ions into YFeO3 or HfO2 can modify FE properties by distortions in the lattice, alterations in electronic structure or formation of new defect complexes [19, 20, 21].
@@
Similarly, the formation of polarons likewise results in the localized state within the band structure. Unlike defects, which are associated with ionic or atomic anomalies, polarons are quasiparticles that form as a result of interaction between charge carriers and lattice vibrations (phonons) [22, 23, 24, 25]. Based on the extent of electron-phonon coupling, polarons are usually classified into large and small polarons, with small polarons showing strong coupling and localized behaviour [25, 26, 27]. Polaron signatures have been detected in several materials such as transition-metal oxides, organic semiconductors, polymers, manganites, hybrid perovskites, cuprites, magnetic semiconductors, and 2D materials, to name a few [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. Charge transport, colossal magnetoresistance, photoemission, surface reactivity, thermoelectricity and (multi)ferroism are a few physical phenomenons where polaron-mediated effects play a crucial role [33, 43, 44, 45, 38, 46, 47, 48, 49].
In FE perovskites such as PbTiO3, theoretical studies have shown that electron polarons become self-trapped and thereby influence the n-type conductivity of the material [50]. Experimental indications that are linked to the material’s optical and transport properties have long suggested the presence and impact of small polarons in BaTiO3. For instance, in BaTiO3 single crystals, the characteristic green luminescence has been linked to small polarons [51, 52]. Furthermore, a proposed explanation for the transport behaviour of charged carriers in BaTiO3 is the concept of a hoping mechanism involving small polarons [53, 54, 55, 56]. Insights into the existence and stability of polarons in BaTiO3 have further been provided by theoretical studies. Computational investigation using Heyd-Scuseria-Ernzerhof (HSE06) functional shows that hole polarons can self-trap in BaTiO3, preferentially localizing on oxygen atoms in both tetragonal and cubic phases [57]. The studies link the experimentally observed low-temperature photoluminescence to the calculated emission peak energy from the recombination of a conduction band electron with a self-trapped hole polaron. Similarly, DFT+U studies also confirm the stability of hole polarons in BaTiO3’s cubic phase [58].
A unique opportunity to explore novel functionalities in materials is presented by understanding the interplay between FE order and small polarons, gaining insight into whether and how FE distortions ($\delta_{Fe}$) can favor or disfavor polaron formation. FE polarons result from such an constructive interplay and are described as polarons stabilized by $\delta_{Fe}$. The presence of FE large polarons in halide perovskites is found to improve charge transport efficiency and enhance performance in optical applications such as solar cells and light-emitting devices [48, 49]. Furthermore, research has demonstrated the role of polarons in the emergence of polarization in SrTiO3 and CaTiO3 [16, 59]. Moreover, ongoing studies are examining multiferroism induced via small polaron dynamics in BaTiO3 [60, 61].
@@
The main aim of this study is to fill this gap by investigating the formation of electron and hole polarons in three different phases of BaTiO3 using hybrid functionals. The stability of polarons in a localized state is compared to its delocalized counterpart across the cubic, tetragonal and rhombohedral polymorphs of BaTiO3 (see Fig. 1). The analysis of the variation of polaron formation energy across the different structures and the reciprocal impact of polarons on the $\delta_{Fe}$, which here is attributed to the relative displacement of Ti and O atoms, is the major component of the current study. Our research elucidate the complex interplay between ferroelectricity and small polarons by comparing the effects of electron and hole polarons on the FE distortive field and polarization in BaTiO3.
II Methods
Density Functional Theory (DFT) stands as a powerful tool in the electronic structures calculations [64, 65]. While this is the case, the limitations of DFT are not unfamiliar, especially when it comes to the study of polarons [66, 67]. For modelling systems where electron localization is pivotal, DFT often leads to inaccuracies due to their inherent self-interaction errors. The widely adopted approach to address this issue is the integration of the Hubbard U term [68]. The description of localized charge carriers are enhanced with DFT$+$U method by adding an on-site coloumbic interaction and thereby correcting the self-interaction errors [69].
[…]
@@ II Methods
An alternative approach to overcome this limitation was to utilize the hybrid functionals [70]. Specifically, we adopted the HSE06 formalism [71]. The fraction of exact Hartree-Fock exchange was set to the standard value of $\alpha$ = 0.25. The first-principle calculations were carried out using the Vienna Ab initio Simulation Package (VASP) in combination with the Projector Augmented Wave (PAW) approach [72, 73, 74, 75]. A plane wave energy cut-off of 320 eV was utilized in all calculations, with 10-5 eV set as the convergence threshold of the electronic self-consistency. The polaronic systems were modelled with a 3x3x3 supercell, containing 135 atoms. The structural relaxation of the ionic positions was continued until the Hellman-Feyman forces were lower than 0.01 eV/Å, and a 2x2x2 Monkhorst-Pack k-point mesh was utilized for the brillouin zone integration. Employing HSE06 for lattice relaxations stabilized the $\delta_{Fe}$ within the ferroelectric systems and concurrently provided lattice parameters and band gaps that more closely agreed to the experimental values[76, 8, 77].
For the formation of electron polarons, an extra electron was introduced into the system and conversely for formation of hole polarons, an electron was removed. Initially, the result was a delocalised solution but by manually breaking the symmetry, it was possible to localize the electron and hole on the preferred site. For electron polarons, six Ti-O bonds around the desired Ti atom was extended, while for hole polarons, two Ti-O bonds next to the chosen O atom was extended. The polaronic formation energy was calculated using the following formula:
@@
Figure 3: The charge density plots due to formation of electron polaron on (a) Cubic BaTiO3 (b) Tetragonal BaTiO3 (c) Rhombohedral BaTiO3. The density of states and the formation of an in-gap state is depicted for (d) Cubic BaTiO3 (e) Tetragonal BaTiO3 (f) Rhombohedral BaTiO3.
@@
Berry-phase approach within the modern theory of polarization was employed to calculate the ferroelectric polarization of both the pristine and polaronic ferroelectric systems [78, 4]. Within the Berry phase approach, the the spontaneous polarization is computed by considering the polarization difference between the ferroelectric phase and a centrosymmetric reference phase, which is the cubic BaTiO3 in the present study. In the context of polaron systems, the considered reference system was a cubic BaTiO3 with a localized charge carrier.
Figure 4: The charge density plots due to formation of hole polaron on (a) Cubic BaTiO3 (b) Tetragonal BaTiO3 (c) Rhombohedral BaTiO3. The density of states and the formation of an in-gap state is depicted for (d) Cubic BaTiO3 (e) Tetragonal BaTiO3 (f) Rhombohedral BaTiO3.
III Results
III.1 Polarons : Electrons and Hole Polarons in BaTiO3
In this section, we inspect the formation of small electron and hole polaron in three polymorphs of BaTiO3 - cubic, tetragonal, and rhombohedral. The distinction among these polymorphs is marked by a varying degree of Ti off-centering within the crystal structure, which shows a progressive increase as we transition from cubic to tetragonal and finally to the rhombohedral, as depicted in Fig.1. The cubic phase of BaTiO3 has a perfectly centered Ti and O atoms , possessing zero off-centering, while the ferroelectric tetragonal and rhombohedral phases are characterized by a Ti off-centering along [001] and [111] directions respectively, leading to $\delta_{Fe}$ in the systems. For the tetragonal phase, we calculate this $\delta_{Fe}$ to be $\sim 0.16$ along the [001] direction, while for the rhombohedral phase, the $\delta_{Fe}$ is computed to be $\sim 0.10$ along the [111] direction. The crystal symmetry and its characteristic FE distortive field play a critical role in the formation and properties of the polaron.
@@ III.1 Polarons : Electrons and Hole Polarons in BaTiO3
Analysis of the charge density plots confirmed the formation of small polarons and the spatial distribution of charge. For electron polarons, a Ti-3d orbital character was exhibited across all three polymorphs, as shown in Fig.3, while an O-2p orbital character was exhibited for hole polarons in all three polymorphs, as shown in Fig.4. For the electron polarons in the cubic and rhombohedral phases, we observed the dxz orbital character, while the preferential occupation of the extra electron in the tetragonal phase was the dxy orbital. For hole polarons, we observed an admixture of py and pz orbital character for cubic phase, a py orbital character for tetragonal phase, while a pz orbital character was observed for the rhombohedral phase of BaTiO3.
@@
Figure 5: The comparison of polaron formation energy for different phases of BaTiO3 for both the electron and hole polaron. The figure depicts the enhanced stability of electron polaron for rhombohedral phase, while the same is not observed in the case of hole polaron, indicating the effect of [111] off-centering of Ti atom in stabilizing the polaron.
@@
Further insights on polaron stability can be obtained from the dependence of EPOL on the crystal symmetry. The stability of electron polarons in BaTiO3 exhibits a clear trend with crystal symmetry, increasing from cubic to tetragonal to rhombohedral phases as shown in Fig. 5. This can be attributed to two reasons: first, the level of antibonding hybridization between the Ti-3d and O-2p orbitals at the conduction band minimum and second, the degree of Ti off-centering. The rhombohedral BaTiO3 exhibits an EPOL of -0.26 eV, which is the lowest of all the three phases considered. The cubic and tetragonal phases show an EPOL of -0.07 eV and -0.11 eV respectively, clearly showing how rhombohedral phase outperforms the other two phases. As mentioned previously, this is attributed to the strong anti-bonding hybridization in this phase, thus allowing maximum energy release upon electron localization on the Ti-3d orbital. A large electronic gain energy (EEL) of -0.49 eV compared to the tetragonal phases’ -0.34 eV validates this. The decrease in $\delta_{Fe}$ of the polaronic octahedron is more significant for the rhombohedral phase than for the tetragonal phase (discussed in detail in the next section), which further validates our observation because the $\delta_{Fe}$ in BaTiO3 is strongly associated with the hybridization between Ti-3d and O-2p. Furthermore, the FE tetragonal and rhombohedral phases, because they already possess off-centering, require less energy to distort compared to the cubic phase. This is clear from the structural energy cost (EST), which is calculated to be 0.27 eV from cubic, while for the tetragonal and rhombohedral phase, this is calculated to be 0.23 eV. These results align with previous findings by Tsunoda et al., highlighting the importance of anti-bonding hybridisation and local displacement of Ti ions along the [111] direction in both rhombohedral BaTiO3 and structurally disordered cubic phases for stabilizing self-trapped polarons. [60]
@@
Interestingly, this trend is not exactly mimicked by hole polarons, as they are already evidently stable in all three polymorphs. Polaron formation energies of -0.20 eV, -0.22 eV, and -0.25 eV are recorded for cubic, tetragonal and rhombohedral BaTiO3, respectively. Although the trend of an increase in stability as one transitions from cubic to rhombohedral is also observed for the hole polaron system, the differences are modest, as shown in Fig.5. Unlike electron polarons, holes are localized on O-2p bonding orbitals, and stabilization involves weakening the bonding O-2p and Ti-3d hybrids. Because the bonding states are already lower in energy and more stable, the hole-polaron formation results in a lower energy gain and smaller variations in the hole-polaron stability. Symmetry also plays a smaller role in the case of hole polarons because the distortion of the O atom, on which the hole polaron forms, from its centrosymmetric position in the ferroelectric phases is computed to be approximately 90% less than that of Ti distortion. The combined effect could possibly explain why a small phase-dependent variation in EPOL was observed in the case of hole polarons. The slight increase in EPOL can be attributed to the pre-distorted TiO6 octahedron in the ferroelectric phases, which could slightly ease the relaxation of the O-site.

### v4

@@ Abstract
In this study, we investigate the formation of electron and hole small polarons in the prototypical ferroelectric material BaTiO3, with a focus on their interaction with ferroelectric distortive fields. To accurately describe the ferroelectric phase in electronically correlated BaTiO3, we employ the HSE06 hybrid density functional, which addresses the limitations of conventional density functional theory (DFT) and Hubbard-corrected DFT+U models, providing a more precise depiction of both ferroelectric and polaronic behaviors. Our analysis spans three structural phases of BaTiO3: cubic, tetragonal, and rhombohedral. We uncover a unique phase-dependent trend in electron polaron stability, which progressively increases across the structural phases, peaking in the rhombohedral phase due to the constructive coupling between the polaron and ferroelectric phonon fields. In contrast, hole polarons exhibit a stability pattern largely unaffected by the phase transitions. Furthermore, we observe that polaron self-trapping significantly alters the local ferroelectric distortive pattern, which propagates to neighboring sites but has a minimal effect on the long-range macroscopic spontaneous polarization. Charge trapping is also associated with localized spin formation, opening new possibilities for enhanced functionalities in multiferroic materials.
@@ I Introduction
Figure 1: (a) Cubic BaTiO3 with zero Ti off-centering (>>393K) (b) Tetragonal BaTiO3 with Ti off-centering in [001] direction (393K - 278K) and (c) Rhombohedral BaTiO3 with Ti off-centering in the [111] direction(<<183K). The orthorhombic phase is in the temperature range of 183-278K but its structure is not shown here because it was not considered in the present study. Detailed investigations of polaronic effects in the orthorhombic phase have been reported[1]. The insets represent the Ti-O bond-lengths in each phase.
@@
Marked by a spontaneous electric polarization (Ps) that can be reversed by an external electric field, ferroelectric (FE) materials are vital in a wide range of technological applications [2, 3, 4, 5]. Capitalizing on this unique switchable polarization property, their applications range from capacitors and transducers to nonvolatile memory devices [6, 7, 8]. Barium titanate (BaTiO3) has been extensively studied, notably for its rich temperature-dependent phase transitions from the high-temperature cubic phase to the tetragonal, orthorhombic, and rhombohedral phases at lower temperatures [9, 10]. This polymorphism opens avenues for investigating the correlation between the electronic structure and lattice distortions, as well as their evolution across different crystal symmetries.
The FE behavior and dielectric properties are altered due to the localized electronic states introduced by the presence of defects, such as oxygen vacancies or cation substitutions. In BaTiO3 and related perovskites such as Pb(ZrxTi1-x)O3, oxygen vacancies act as electron donors, potentially creating defect dipoles that pin domain walls, thereby affecting polarization switching [11, 12, 13, 14]. Similarly, for BaTiO3 and SrTiO3, titanium vacancies or antisite defects were shown to reduce or modify the ferroelectricity [15, 16, 17, 18, 19]. Such defect-induced localization plays a pivotal role in tuning the electronic and ferroic properties of materials, thereby affecting both the static polarization and dynamic processes, such as charge transport.
@@
Similarly, polaron formation results in a localized state within the band structure. Unlike defects, which are associated with ionic or atomic anomalies, polarons are quasiparticles that form as a result of the interaction between charge carriers and lattice vibrations (phonons) [20, 21, 22, 23]. Polarons are usually classified into large and small polarons based on the extent of electron-phonon coupling, with small polarons showing strong coupling and localized behavior [20, 21, 22, 23, 23, 24, 25]. Polaron signatures have been detected in several materials, such as transition metal oxides, organic semiconductors, polymers, manganites, hybrid perovskites, cuprites, magnetic semiconductors, and 2D materials [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. Charge transport, colossal magnetoresistance, photoemission, surface reactivity, thermoelectricity, and (multi)ferroism are a few physical phenomena in which polaron-mediated effects play a crucial role [31, 42, 43, 44, 37, 45, 46, 47, 48, 49]..
Theoretical and experimental studies have provided ample evidence for the presence of polarons in BaTiO3. Experimental indications linked to the optical and transport properties of materials have long suggested the presence and impact of small polarons in BaTiO3. For instance, in BaTiO3 single crystals, the characteristic green luminescence has been linked to the presence of small polarons, and the transport behavior has been explained through hopping mechanisms involving these quasiparticles [50, 51, 52, 53, 54, 55]. Computational investigations using hybrid functionals have shown that hole polarons can self-trap on oxygen atoms in BaTiO3 in both the tetragonal and cubic phases, consistent with the experimentally observed low-temperature photoluminescence [56]. DFT+U studies also confirmed the stability of hole polarons in BaTiO3’s cubic phase [57]. Understanding the interplay between the FE order and small polarons provides a unique opportunity to explore novel functionalities of this material. Gaining insights into whether and how FE distortions ($\delta_{FE}$) in BaTiO3, affect polaron formation is essential for understanding the stability and behavior of polarons. In BaTiO3, these distortions correspond to an off-centering of the Ti atom from the center of the TiO6 octahedron, along with a small movement of the O atom, leading to a Ps. FE polarons result from such a constructive interplay and are described as polarons stabilized by $\delta_{FE}$. Such polarons have been shown to enhance the charge transport efficiency and optical performance of related materials, such as halide perovskites [47, 48]. Moreover, ongoing studies are examining multiferroism induced by small polarons in BaTiO3 and similar perovskites, highlighting the broader significance of polaron-ferroelectric coupling in these oxides [58, 1, 59].
@@
The main aim of this study is to fill this gap by investigating the formation of electron and hole polarons in three different phases of BaTiO3 using hybrid functionals. The main focus of the current study is the analysis of the variation in polaron stability across different structures and the reciprocal impact of polarons on $\delta_{FE}$. By elucidating this complex interplay between ferroelectricity and small polarons across different phases and charge carrier types, our research advances the understanding of the complex interrelations between the FE order and charge localization in BaTiO3.
II Material background and computational approach
As mentioned earlier, BaTiO3 is a classic ferroelectric material that exhibits multiple structural phases with distinct polarization behaviors, as illustrated in Figure 1. The cubic phase of BaTiO3 has perfectly centered Ti and O atoms, possessing zero off-centering from the TiO6 octahedron, while the ferroelectric tetragonal and rhombohedral phases are characterized by Ti off-centering along the [001] and [111] directions, respectively, leading to $\delta_{FE}$ in the systems. For the tetragonal phase, we calculated $\delta_{FE}$ to be $\sim 0.16$ Å along the [001] direction, while for the rhombohedral phase, $\delta_{FE}$ was computed to be $\sim 0.10$ Å along the [111] direction.
The bond lengths vary significantly among the phases according to these distortions. In the cubic phase, all the bonds are equal in length. As shown in Figure 1(b), the in-plane bonds of tetragonal BaTiO3 are equal in length, whereas the out-of-plane bonds are of two different lengths, one short and the other long. The rhombohedral phase of BaTiO3 also has two distinct bond types according to their length, three of which are longer and the other three are shorter, as shown in Figure 1(c). The variation in bond lengths corresponds to the changes in the orbital hybridization between the Ti-3dd and O-2pp states, which plays a key role in stabilizing the ferroelectric phase in BaTiO3 through enhanced covalent interactions. The modulation of this hybridization is fundamental to understanding the behavior of polarons and the FE stability of the material.
While the orthorhombic phase of BaTiO3 also exhibits ferroelectricity, the present study excludes this because its polaron behavior has been extensively studied previously[1]. Moreover, our focus on the tetragonal and rhombohedral phases is also due to their representation of uniaxial and triaxial polarization symmetries, respectively, which are more relevant to our study, which focuses on polarization-directional effects.
From a computational perspective, Density Functional Theory (DFT) is a powerful tool for electronic structure calculations [62, 63]; however, the limitations of DFT are not unfamiliar, especially when it comes to the study of polarons [64, 65]. For modelling systems where electron localization is pivotal, DFT often leads to inaccuracies due to inherent self-interaction errors. The widely adopted approach to address this issue is the integration of the Hubbard U term into DFT calculations, which corrects these errors by accounting for the onsite coloumbic interaction [66, 67]. However, DFT+U poses challenges to the current study. The stabilization of $\delta_{FE}$ within FE systems is vital to this study. However, a linear reduction in $\delta_{FE}$ was observed for the FE phases of BaTiO3 as the U value increased from 0 to 5 eV and beyond, resulting in a paraelectric cubic structure [60, 61]. This observation was evident for both the tetragonal and rhombohedral phases during geometrical optimization. The stronger localization of electrons in the d-orbitals of the Ti atoms arises with the increase in the U value, consequentially diminishing the covalency due to a suppressed hybridization between the O-2pp and Ti-3dd orbitals. Figure 2 illustrates the phonon dispersion curve of cubic BaTiO3 and clearly demonstrates that the PBE+U method does not show any imaginary phonon modes. The absence of these modes indicates that cubic BaTiO3 does not exhibit soft phonon modes associated with FE instability. The inability of DFT+U to stabilize $\delta_{FE}$ has already been reported and studied[60, 61]. An alternative approach to overcome this limitation is to utilize hybrid functionals, thereby providing a balanced treatment to study polaronic phenomena while preserving the ferroelectricity in BaTiO3’s FE phases [68].
@@ II Material background and computational approach
Building on this understanding of the structural characteristics of BaTiO3 and the validated computational approach, we detail the specific methodologies employed in our study to investigate polaron behavior within these ferroelectric phases.
III Methods
We adopted the Heyd-Scuseria-Ernzerhof (HSE06) formalism [69]. The fraction of exact Hartree-Fock exchange was set to the standard value of $\alpha$ = 0.25. The first-principle calculations were carried out using the Vienna Ab initio Simulation Package (VASP) in combination with the Projector Augmented Wave (PAW) approach [70, 71, 72, 73].The pseudopotentials used in our calculations are Basv (5s25p66s2), Ti (3d34s1), and Os(2s22p4). A plane wave energy cut-off of 320 eV was utilized in all calculations, with 10-5 eV set as the convergence threshold of the electronic self-consistency. The polaronic systems were modelled with a 3x3x3 supercell, containing 135 atoms. The structural relaxation of the ionic positions was continued until the Hellman-Feyman forces were lower than 0.01 eV/ Å, and a 2x2x2 Monkhorst-Pack k-point mesh was utilized for the brillouin zone integration. Employing HSE06 for lattice relaxations stabilized the $\delta_{FE}$ within the ferroelectric systems and concurrently provided lattice parameters and band gaps that more closely agreed to the experimental values[74, 9, 75].
For the formation of electron polarons, an extra electron was introduced into the system and conversely for formation of hole polarons, an electron was removed. Simultaneously, a compensating background charge of opposite sign is included to keep the system neutral. This method can mimic the experimental conditions in which polarons are formed through photoexcitation rather than explicit defects, providing a computationally efficient and physically relevant approach for studying polaron formation and behavior[76, 77, 78, 79]. Initially, the result was a delocalised solution but by manually breaking the symmetry, it was possible to localize the electron and hole on the preferred site. For electron polarons, six Ti-O bonds around the desired Ti atom were extended, while for hole polarons, two Ti-O bonds next to the chosen O atom were extended. The polaronic formation energy was calculated using the following formula:
@@ III Methods
$\begin{split}E_{ST}=E_{dist}^{deloc}-E_{unif}^{deloc}\end{split}$ (2)
$\begin{split}E_{EL}=E_{dist}^{loc}-E_{dist}^{deloc}\end{split}$ (3)
[…]
@@
Figure 3: The charge density plots presented in the top panel illustrate electron polaron isosurfaces in (a) Cubic BaTiO3, (b) Tetragonal BaTiO3, and (c) Rhombohedral BaTiO3. Meanwhile, the bottom panel depicts hole polaron isosurfaces in (d) Cubic BaTiO3, (e) Tetragonal BaTiO3, and (f) Rhombohedral BaTiO3.
@@
Berry-phase approach within the modern theory of polarization was employed to calculate the ferroelectric polarization of both the pristine and polaronic ferroelectric systems [80, 5]. Within the Berry-phase approach, the the Ps is computed by considering the polarization difference between the ferroelectric phase and a centrosymmetric reference phase, which is the cubic BaTiO3 in the present study. In the context of polaron systems, the considered reference system was a cubic BaTiO3 with a localized charge carrier.
IV Results
With our results, we attempt to unravel how electron and hole polarons couple to ferroelectricity in BaTiO3, first by examining the role of $\delta_{FE}$ in the stability of the polarons (Section A), and then by analyzing the impact of these localized entities on the local ferroelectric order in the ferroelectric phases of the material (electron polarons: Section B; hole polarons: Section C).
IV.1 Polarons : Electrons and Hole Polarons in BaTiO3
@@
To acheive our goals, we inspect the formation of small electron and hole polaron in three polymorphs of BaTiO3 - cubic, tetragonal, and rhombohedral. Analysis of the charge density plots confirmed the formation of small polarons and the spatial distribution of charge. As shown in Figure 3, a Ti-3dd orbital character was observed across all three polymorphs for electron polarons, while an O-2pp orbital character was exhibited for hole polarons in all three polymorphs. For the electron polarons in the cubic phase, we observed an admixture of dxy and dyz orbital character, while the preferential occupation of the extra electron in the tetragonal phase was the dxy orbital. For the rhombohedral phase, electron polaron occupies the dxz orbital. For hole polarons, we observed an admixture of py and pz orbital character for cubic phase, a py orbital character for tetragonal phase, while a px orbital character was observed for the rhombohedral phase of BaTiO3. The density of states (DOS) plots for electron and hole polarons in different phases of BaTiO3 are provided in the Figure S1 of supplementary materials (SM)[81].
@@
Figure 4: The comparison of polaron formation energy for different phases of BaTiO3 for both the electron and hole polaron. The figure depicts the enhanced stability of electron polaron for rhombohedral phase, while the same is not observed in the case of hole polaron, indicating the effect of [111] off-centering of Ti atom in stabilizing the polaron. The polarization direction is indicated within square brackets and is not shown in the case of the cubic phase because it is non-polar.
@@
Further insights on polaron stability can be obtained from the dependence of EPOL on the crystal symmetry. The stability of electron polarons in BaTiO3 exhibits a clear trend with crystal symmetry, increasing from cubic to tetragonal to rhombohedral phases as shown in Figure 4. This can be attributed to two reasons: first, the level of antibonding hybridization between the Ti-3dd and O-2pp orbitals at the conduction band minimum and second, the degree of Ti off-centering. The rhombohedral BaTiO3 exhibits an EPOL of -0.26 eV, which is the lowest of all the three phases considered. The cubic and tetragonal phases show an EPOL of -0.07 eV and -0.11 eV respectively, clearly showing how rhombohedral phase outperforms the other two phases. As mentioned previously, this is attributed to the strong anti-bonding hybridization in this phase, thus allowing maximum energy release upon electron localization on the Ti-3dd orbital. A large electronic gain energy (EEL) of -0.49 eV compared to the tetragonal phase’s -0.34 eV validates this. The decrease in $\delta_{FE}$ of the polaronic octahedron is more significant for the rhombohedral phase than for the tetragonal phase (discussed in detail in the next section), which further validates our observation because the $\delta_{FE}$ in BaTiO3 is strongly associated with the hybridization between Ti-3dd and O-2pp. Furthermore, the FE tetragonal and rhombohedral phases, because they already possess off-centering, require less energy to distort compared to the cubic phase. This is clear from the structural energy cost (EST), which is calculated to be 0.27 eV from cubic, while for the tetragonal and rhombohedral phase, this is calculated to be 0.23 eV. These results align with previous findings by Tsunoda et al., highlighting the importance of anti-bonding hybridisation and local displacement of Ti ions along the [111] direction in both rhombohedral BaTiO3 and structurally disordered cubic phases for stabilizing self-trapped polarons. [58]
@@
Interestingly, this trend is not exactly mimicked by hole polarons, as they are already evidently stable in all three polymorphs. Polaron formation energies of -0.20 eV, -0.22 eV, and -0.25 eV are recorded for cubic, tetragonal and rhombohedral BaTiO3, respectively. Although the trend of an increase in stability as one transitions from cubic to rhombohedral is also observed for the hole polaron system, the differences are modest, as shown in Figure 4. Unlike electron polarons, holes are localized on O-2pp bonding orbitals, and stabilization involves weakening the bonding O-2pp and Ti-3dd hybrids. Because the bonding states are already lower in energy and more stable, the hole-polaron formation results in a lower energy gain and smaller variations in the hole-polaron stability. Symmetry also plays a smaller role in the case of hole polarons because the distortion of the O atom, on which the hole polaron forms, from its centrosymmetric position in the ferroelectric phases is computed to be approximately 90% less than that of Ti distortion. The combined effect could possibly explain why a small phase-dependent variation in EPOL was observed in the case of hole polarons. The slight increase in EPOL can be attributed to the pre-distorted TiO6 octahedron in the ferroelectric phases, which could slightly ease the relaxation of the O-site.